# UAV-Rain1k: A Benchmark for Raindrop Removal from UAV Aerial Imagery

### 08 Feb 2024 v1

@@ Abstract
Raindrops adhering to the lens of UAVs can obstruct visibility of the background scene and degrade image quality. Despite recent progress in image deraining methods and datasets, there is a lack of focus on raindrop removal from UAV aerial imagery due to the unique challenges posed by varying angles and rapid movement during drone flight. To fill the gap in this research, we first construct a new benchmark dataset for removing raindrops from UAV images, called UAV-Rain1k. In this letter, we provide a dataset generation pipeline, which includes modeling raindrop shapes using Blender, collecting background images from various UAV angles, random sampling of rain masks and etc. Based on the proposed benchmark, we further present a comprehensive evaluation of existing representative image deraining algorithms, and reveal future research opportunities worth exploring. The proposed dataset will be publicly available at https://github.com/cschenxiang/UAV-Rain1k.
Index Terms:
Raindrop removal, image deraining, UAV aerial image, benchmark, dataset.
I Introduction
As unmanned aerial vehicles (UAVs) are increasingly employed for tasks such as surveillance, mapping, and environmental monitoring, the quality of captured imagery becomes paramount. On rainy days, raindrops adhering to the lens of UAVs can significantly degrade image quality, hindering visibility and affecting the accuracy of downstream image processing tasks. Thus, developing effective raindrop removal techniques tailored to UAV imagery is essential for enhancing the reliability and usability of UAV-based applications [1].
@@ I Introduction
In recent years, notable progress has been made in image deraining, propelled by the advent of effective image priors and deep learning networks [2]. To better address the problem of removing raindrops, Qian et al. [3] constructed the first raindrop removal dataset, comprising 1,119 pairs of images featuring raindrops against varied backgrounds. These images are captured using a camera equipped with two aligned pieces of glass, one sprayed with water while the other remained clean. Later, Quan et al. [4] presented a mixed rain dataset, namely RainDS, which contains synthetic and real-world raindrops created by manually mimicking rainfall with a sprinkler.
@@
The rest of this letter is structured as follows. In Section II, we provide a detailed description of the pipeline used to construct the dataset. We analyze the performance of existing image deraining algorithms on our benchmark in Section III. Finally, the concluding remarks will be given in Section IV.
Figure 2: Illustration of the dataset generation pipeline.
II Dataset Construction
@@ II-A Raindrop Generation
Inspired by Grag [8], we model the generation of raindrop layers as a motion blur process. Mathematically, this motion process can be represented based on the oscillation characteristics of falling raindrops:
$r=r_{0}+p(\theta).$ (1)
The shape of a water droplet at any given time is represented by $r$, where $r$ is the distance from a point on the droplet’s surface to its center. As the droplet falls, it undergoes rapid shape distortion over time, with $r_{0}$ as the undeformed radius and $\theta$ as the rotation angle during the droplet’s descent. The function $p(\theta)$ describes the cumulative influence of modal shapes and amplitudes on the angle $\theta$ over time.
II-B Background Collection
The quality of the clear background image, also known as the ground-truth image, is equally crucial when constructing a paired dataset [9, 2]. These images should have higher resolution and rich object details. Unfortunately, existing datasets have placed significant emphasis on rain simulation but have overlooked the need for high-quality background images. Regrettably, the realistic rainy weather images used in previous synthetic datasets were consistently captured from a flat angle on the ground. In reality, there is a significant disparity between the lighting conditions in the air and those on the ground, especially in aerial imagery. Aerial images predominantly feature ground structures as the primary compositional elements in the background, which aligns with human visual perception of rainy scenes. Obtaining rain-free background images that accurately represent the human perception of real rainy conditions is a challenging task, particularly when searching for such images on the Internet. These images also suffer from various issues, including compression artifacts, watermarks, low resolution, defocused blurriness, absence of objects, and etc. These challenges can present difficulties for advanced vision-based remote sensing applications.
Our collection of background images encompasses a wide range of scenes, including urban, natural, and rural environments. Furthermore, we deliberately sought variations in aerial shooting angles during the photography process to capture different scene effects in real-world scenarios. Consequently, the complexity of these scenes and backgrounds presents both challenges and value to our subset of real-world data.
@@ II-B Background Collection
II-C Image Composition
In [3], a method is employed where a piece of glass with attached water droplets is inserted in front of the drone camera lens to simulate the presence of raindrops. However, during the process of capturing aerial rain images, due to the complexity and specificity of capturing scenes with moving perspectives, we are unable to obtain perfectly consistent rainy and non-rainy images through aerial images. This method also provides some insights for verifying the authenticity of synthetic datasets. We collect real aerial images captured during rainy weather and gather drone footage of raining scenes at different angles and motion states as a reference baseline for synthetic images. This ensures that the collected images exhibit variations in raindrop density, size, and positions, allowing us to filter out low-quality and misleading synthetic data.
Our goal is to achieve visual realism and consistency in synthetic rainy images, with the aim of minimizing the domain gap between synthetic and real images. To achieve this, we do not rely on simple copy-and-paste methods. Instead, we employ a random sampling approach for raindrop masks to ensure that each synthesized rainy image is unique. Subsequently, we use Photoshop for rendering to seamlessly integrate the raindrop masks with the background images. As a result, compared to other dataset generation methods, our proposed datasets tend to be more complex and diverse, enhancing their ability to generalize to various scenarios.
@@ II-C Image Composition
II-D Benchmark Statistics
As a result, we propose a new benchmark for raindrop removal from UAV aerial imagery, called UAV-Rain1k. In total, the training and testing set of the UAV-Rain1k contains 800 and 220 synthetic images, respectively. The average resolution of all images is $1500\times 1000$. Our dataset comprises four rain density labels (e.g., light, moderate, heavy, and irregular). Furthermore, we categorize the background images according to aerial shooting angles and aerial scenes. Consequently, our dataset demonstrates greater diversity and a larger scale in terms of both rain conditions and ground truth (GT) images. Our emphasis goes beyond just the sequence number and total frames of real rain. We also prioritize the diversity of real rain patterns and the authenticity of background images. This focus has a significant impact on the model’s ability to generalize to real rain scenarios. Figure 5 provides some visual examples.
II-E Subjective Assessment
@@
III Algorithm Benchmarking
TABLE I: Quantitative comparisons on the UAV-Rain1k benchmark dataset. “#FLOPs” and “#Params” represent FLOPs (in G) and the number of trainable parameters (in M), respectively.
@@ III Algorithm Benchmarking
PSNR $\uparrow$ 16.74 16.62 22.36 22.43 25.15 22.10 24.20
SSIM $\uparrow$ 0.7142 0.7088 0.8686 0.8529 0.9101 0.8861 0.8994
Params $\downarrow$ - - 3.17 3.04 26.12 16.41 33.65
FLOPs $\downarrow$ - - 21.2 89.3 174.7 61.9 242.9
III-A Quantitative Evaluation
Table I presents the quantitative performance evaluation results of various algorithms on the UAV-Rain1k dataset. Here, we utilize two widely referenced image quality assessment metrics, PSNR (Peak Signal-to-Noise Ratio) [16] and SSIM (Structural Similarity Index) [17], to measure the restoration quality of different methods. Additionally, we assess the model efficiency of each approach by considering parameters and FLOPs (Floating Point Operations per Second).
@@ III-A Quantitative Evaluation
Figure 6: Visual quality comparison on the UAV-Rain1k dataset. Zooming in the figures offers a better view at the deraining capability.
III-B Qualitative Evaluation
Figure 6 illustrates the visual comparison results of different baseline algorithms on our proposed benchmark. In comparison to CNN-based methods (RCDNet [11] and SPDNet [12]), Transformer-based approaches exhibit superior capability in removing undesired raindrop and recovering clearer visual outcomes, owing to the evident advantage of Transformers in modeling non-local information. Additionally, we observe grid artifacts in the restoration results of IDT [14] and DRSformer [15], indicating their limited capacity in handling high-resolution UAV imagery. Furthermore, nearly all methods exhibit poor performance in local detail regions, as they tend to remove some image details while removing raindrop, suggesting that there are still rooms for improvement.
III-C Application-Based Evaluation
@@
IV Conclusion
In this letter, we have proposed a new benchmark dataset for raindrop removal from UAV aerial imagery. Based on the proposed new benchmark, we further experimentally validated representative image deraining algorithms in three aspects: quantitative evaluation, qualitative evaluation, and application-based evaluation. The experimental results shed light on the limitations of existing methods and point towards promising future directions. This letter not only pioneers a new research direction in UAV image processing but also advocates for the community to propose more effective algorithms.

### 08 Apr 2024 v2

@@ Abstract
Raindrops adhering to the lens of UAVs can obstruct visibility of the background scene and degrade image quality. Despite recent progress in image deraining methods and datasets, there is a lack of focus on raindrop removal from UAV aerial imagery due to the unique challenges posed by varying angles and rapid movement during drone flight. To fill the gap in this research, we first construct a new benchmark dataset for removing raindrops from UAV images, called UAV-Rain1k. In this paper, we provide a dataset generation pipeline, which includes modeling raindrop shapes using Blender, collecting background images from various UAV angles, random sampling of rain masks and etc. Based on the proposed benchmark, we further present a comprehensive evaluation of existing representative image deraining algorithms, and reveal future research opportunities worth exploring. The proposed dataset is publicly available at https://github.com/cschenxiang/UAV-Rain1k.
1 Introduction
As unmanned aerial vehicles (UAVs) are increasingly employed for tasks such as surveillance, mapping, and environmental monitoring, the quality of captured imagery becomes paramount. Aerial images are susceptible to degradation in clarity, color distortion, and loss of information due to adverse weather conditions and other factors during the acquisition process [20]. For example, during rainy weather, raindrops adhering to the lens of UAVs can significantly degrade image quality, hindering visibility and affecting the accuracy of downstream image processing tasks. Thus, developing effective raindrop removal techniques tailored to UAV imagery is essential for enhancing the reliability and usability of UAV-based applications [2, 21, 18].
In recent years, notable progress has been made in image deraining, propelled by the advent of effective image priors and deep learning networks [4, 1]. To better address the problem of removing raindrops, Qian et al. [12] constructed the first raindrop removal dataset, comprising 1,119 pairs of images featuring raindrops against varied backgrounds. These images are captured using a camera equipped with two aligned pieces of glass, one sprayed with water while the other remained clean. Later, Quan et al. [13] presented a mixed rain dataset, namely RainDS, which contains synthetic and real-world raindrops created by manually mimicking rainfall with a sprinkler.
@@ 1 Introduction
The rest of this paper is structured as follows. In Section 2, we provide a detailed description of the pipeline used to construct the dataset. We analyze the performance of existing algorithms on our benchmark in Section 3. Finally, the concluding remarks will be given in Section 4.
2 Dataset Construction
@@ 2.1 Raindrop Generation
Inspired by Garg [5], we model the generation of raindrop layers as a motion blur process. The instantaneous shape of a water droplet at time $t$ is represented by the function $r[t,\vartheta,\phi]$, where $r$ is the distance from a point on the droplet’s surface to its center, $\theta$ is defined as the angle opposite to the direction of the droplet’s fall at that point, and $\phi$ is defined as the angle between that point and the projection of the line of sight onto any plane perpendicular to the direction of fall. As raindrops fall, the influence of aerodynamics and surface tension acting on the raindrop causes rapid shape distortion over time. Mathematically, this motion process can be represented based on the oscillatory characteristics of the falling raindrop:
$r[t,\vartheta,\phi]={r}_{0}(1+\displaystyle\sum_{m}\cos(m\phi){p}_{m}(\theta)),$ (1)
where ${r}_{0}$ is the undeformed radius, and the factor $\cos(m\phi)$ depends on the droplet size ${r}_{0}$. The function ${p}_{m}(\theta)$ describes the time dependent variation of modal shape and amplitude with respect to $\theta$. The droplet’s shape at any given moment is determined by the cumulative impact of falling time on both modal shape and amplitude.
2.2 Background Collection
The quality of the clear background image, also known as the ground-truth image, is equally crucial when constructing a paired dataset [6, 4, 10]. These images should have higher resolution and rich object details. Unfortunately, existing datasets have placed significant emphasis on rain simulation but have overlooked the need for high-quality background images. Regrettably, the realistic rainy weather images used in previous synthetic datasets were consistently captured from a flat angle on the ground. In reality, there is a significant disparity between the lighting conditions in the air and those on the ground, especially in aerial imagery. Aerial images predominantly feature ground structures as the primary compositional elements in the background, which aligns with human visual perception of rainy scenes. Obtaining rain-free background images that accurately represent the human perception of real rainy conditions is a challenging task, particularly when searching for such images on the Internet. These images also suffer from various issues, including compression artifacts, watermarks, low resolution, defocused blurriness, absence of objects, and etc. These challenges can present difficulties for advanced vision-based remote sensing applications.
@@ 2.2 Background Collection
Our collection of background images encompasses a wide range of scenes, including urban, natural, and rural environments. Furthermore, we deliberately sought variations in aerial shooting angles during the photography process to capture different scene effects in real scenarios. Consequently, the complexity of these scenes and backgrounds presents both challenges and value to our subset of real data.
@@
Table 1: Quantitative comparisons on the UAV-Rain1k benchmark dataset. “#FLOPs” and “#Params” represent FLOPs (in G) and the number of trainable parameters (in M), respectively.
@@
PSNR 16.74 16.62 22.36 22.43 25.15 22.10 24.20
SSIM 0.7142 0.7088 0.8686 0.8529 0.9101 0.8861 0.8994
#Params - - 3.17 3.04 26.12 16.41 33.65
#FLOPs - - 21.2 89.3 174.7 61.9 242.9
2.3 Image Composition
In [12], a method is employed where a piece of glass with attached water droplets is inserted in front of the drone camera lens to simulate the presence of raindrops. However, during the process of capturing aerial rain images, due to the complexity and specificity of capturing scenes with moving perspectives, we are unable to obtain perfectly consistent rainy and rain-free images through aerial images. This method also provides some insights for verifying the authenticity of synthetic datasets. We collect real aerial images captured during rainy weather and gather drone footage of raining scenes at different angles and motion states as a reference baseline for synthetic images. This ensures that the collected images exhibit variations in raindrop density, size, and positions, allowing us to filter out low-quality and misleading synthetic data.
Our goal is to achieve visual realism and consistency in synthetic rainy images, with the aim of minimizing the domain gap between synthetic and real images. To achieve this, we do not rely on simple copy-and-paste methods. Instead, we employ a random sampling approach for raindrop masks to ensure that each synthesized rainy image is unique. To ensure the synthesized images exhibit visual realism and reduce the domain gap between synthesized and real images, inspired by Qian et al. [12], we model the blurred or occluded effects $D$ of raindrops in dispersed small-sized locally coherent regions, combined with a clean background image $B$, to form the degraded raindrop image ${R}_{d}$. Mathematically, the synthesis process can be expressed as follows:
${R}_{d}=(1-M)\bigodot B+D,$ (2)
where $D$ represents the occlusion or blurring effects caused by raindrops, and $B$ is a binary mask. If $M(x)=1$, then the pixel $x$ in the mask belongs to the raindrop region, otherwise, it is part of the background image. As a result, we seamlessly integrate the raindrop mask with the background image. Consequently, compared to other dataset generation methods, our proposed dataset tends to be more complex and diverse, enhancing its ability to generalize across various scenarios.
2.4 Benchmark Statistics
As a result, we propose a new benchmark for raindrop removal from UAV aerial imagery, called UAV-Rain1k. In total, the training and testing set of the UAV-Rain1k contains 800 and 220 synthetic images, respectively. The average resolution of all images is $1500\times 1000$. Our dataset comprises four rain density labels (e.g., light, moderate, heavy, and irregular). In Figure 3, the inner ring displays labels for each rain density, corresponding to a specific quantity of synthesized images. Furthermore, we categorize the background images according to aerial shooting angles and aerial scenes. Consequently, our dataset demonstrates greater diversity and a larger scale in terms of both rain conditions and ground truth (GT) images. Our emphasis goes beyond just the sequence number and total frames of real rain. We also prioritize the diversity of real rain patterns and the authenticity of background images. This focus has a significant impact on the model’s ability to generalize to real rain scenarios.
Based on the aforementioned work, we have summarized three innovative aspects of the UAV-Rain1k dataset:
  • A more comprehensive and reliable background collection effort. Encompassing a diverse range of scenes, while simultaneously considering the complexity and uniqueness of capturing scenes from dynamic perspectives.
  • More authentic raindrop shapes. We model the generation of the raindrop layer as a motion blur process and employed an open source 3D rendering engine for the depiction of raindrops, achieving higher quality and more realistic raindrop masks.
  • A more natural synthesis effect. Integrating authentic and diverse raindrop masks seamlessly with background images covering various scenes to achieve a more realistic visual effect. Figure 5 provides some visual examples.
Figure 6: Visual quality comparison on the UAV-Rain1k dataset (horizontal angle). Zooming in the figures offers a better view at the deraining capability.
Figure 7: Visual quality comparison on the UAV-Rain1k dataset (downward angle). Zooming in the figures offers a better view at the deraining capability.
2.5 Subjective Assessment
3 Algorithm Benchmarking
3.1 Quantitative Evaluation
Table 1 presents the quantitative performance evaluation results of various algorithms on the UAV-Rain1k dataset. Here, we utilize two widely referenced image quality assessment metrics, PSNR (Peak Signal-to-Noise Ratio) [8] and SSIM (Structural Similarity Index) [17], to measure the restoration quality of different methods. Additionally, we assess the model efficiency of each approach by considering parameters and FLOPs (Floating Point Operations per Second).
@@ 3.1 Quantitative Evaluation
3.2 Qualitative Evaluation
Figure 6 and Figure 7 illustrates the visual comparison results of different baseline algorithms on our proposed benchmark (In Figure 6, the effect corresponds to a composite image taken from a horizontal angle, while Figure 7 represents a composite image captured from an downward angle). In comparison to CNN-based methods (RCDNet [16] and SPDNet [22]), Transformer-based approaches exhibit superior capability in removing undesired raindrop and recovering clearer visual outcomes, it achieves remarkable visual effects from any aerial perspective, eliminating raindrops and obscuring effects without rain residues and pseudo artifacts. owing to the evident advantage of Transformers in modeling non-local information. Additionally, we observe grid artifacts in the restoration results of IDT [19] and DRSformer [3], indicating their limited capacity in handling high-resolution UAV imagery. Furthermore, nearly all methods exhibit poor performance in local detail regions, as they tend to remove some image details while removing raindrop, suggesting that there are still rooms for improvement.
3.3 Application-Based Evaluation
@@
4 Conclusion
In this paper, we have proposed a new benchmark dataset for raindrop removal from UAV aerial imagery. We have provided a detailed overview of the synthesis process for the UAVRain1k dataset, covering background collection, raindrop modeling, image synthesis, results demonstration, and subjective evaluation. Based on the proposed new benchmark, we further experimentally validated representative image deraining algorithms in three aspects: quantitative evaluation, qualitative evaluation, and application-based evaluation. The experimental results shed light on the limitations of existing methods and point towards promising future directions. This paper not only pioneers a new research direction in UAV image processing but also advocates for the community to propose more effective algorithms.